# Faithful global convergence for the rescaled Consensus–Based Optimization

### v1

@@ 2. Useful estimates
[…]
@@
    $f:~{}{\mathbb{R}}^{d}\to{\mathbb{R}}_{+}$ is bounded from below by $\underline{f}:=\min f>0$, and there exists some constant $L_{f}>0$ such that
@@
    $f-\underline{f}\leq c_{u}(|x|^{2}+1),\quad\forall x\in{\mathbb{R}}^{d}\;\mbox{and }\;c_{\ell}|x|^{2}\leq f-\underline{f}\mbox{ for }|x|\geq M\,.$
As we have mentioned earlier in the introduction, our main ingredients are moments bounds and a Harnack inequality satisfied by the invariant probability measure. These are the object of the current section.
2.1. Moments bounds
The mean–field term $\mathfrak{m}_{\alpha}(\mu)$ plays a crucial role in the behavior of the dynamics. Its properties are essential for our analysis to be carried out. Thus we shall recall some estimates on $\mathfrak{m}_{\alpha}(\mu)$ that we borrow from [carrillo2018analytical, Lemma 3.3].
@@ 2.1. Moments bounds
Suppose that $f(\cdot)$ satisfies Assumption 2.1, then it holds that
@@
$b_{1}=M^{2}+b_{2}\quad\text{ and }\quad b_{2}=2\,\frac{c_{u}}{c_{\ell}}\left(1+\frac{1}{\alpha\,M^{2}\,c_{\ell}}\right).$
@@
In fact, as it has been already observed in [carrillo2018analytical, Remark .1], the constant $b_{2}$ can be chosen to be independent of $\alpha$ for $\alpha\geq 1$.
@@
Let $f$ satisfy Assumption 2.1 and $\rho_{0}\in\mathscr{P}_{p}({\mathbb{R}}^{d})$ for any $p\geq 2$, then both the dynamics (1.5) and (1.4) have unique strong solutions and they satisfy that
[…]
@@
The proof of well–posedness for the particle system (1.5) follows the same reasoning as in [carrillo2018analytical, Theorem 2.1] or [gerber2023mean, Theorem 2.2], relying on the fact that the coefficients in (1.5) are locally Lipschitz. For the mean-field dynamics (1.4), well-posedness can be established using the Leray-Schauder fixed-point theorem, as demonstrated in [carrillo2018analytical, Theorem 3.1] or [gerber2023mean, Theorem 2.2]. Finally, the moment estimates in (2.2) are derived from [carrillo2018analytical, Lemma 3.4] or [gerber2023mean, Lemma 3.5] by using the Burkholder–Davis–Gundy inequality. ∎
@@
Let $f(\cdot)$ satisfy assumption 2.1 and it satisfies $\rho_{0}\in\mathscr{P}_{4}({\mathbb{R}}^{d})$. Consider $({\overline{X}}_{t})_{t\geq 0}$ with $\rho_{0}$-distributed initial data satisfying (1.4). Then for sufficiently small $\kappa$ there exists some constant $C_{2}$ depending only $\lambda,\sigma,d,\delta,\kappa$ and $f$ such that it holds
(2.3) $\sup_{t\geq 0}{\mathbb{E}}[|{\overline{X}}_{t}|^{2}]\leq\mathbb{E}\left[|{\overline{X}}_{0}|^{2}\right]+C_{2}\,.$
@@
Let $h(x):=\frac{1}{2}|x|^{2}$, for $x\in\mathbb{R}^{d}$. Then applying Itô’s formula leads to
${\rm d}h({\overline{X}}_{t})$ =$=\bigg{\{}-\lambda\nabla h({\overline{X}}_{t})\cdot\big{(}{\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t})\big{)}$
@@
This is true in particular because of the well-posedness of the process ${\overline{X}}_{\cdot}$ according to Lemma 2.4 with $X_{0}^{i}\in\mathscr{P}_{4}({\mathbb{R}}^{d})$, and its finite moments which guarantee that $\mathbb{E}\left[\int_{0}^{t}|G(s,{\overline{X}}_{s})|^{2}{\rm d}s\right]\leq C\,{\mathbb{E}}\left[\int_{0}^{t}|{\overline{X}}_{s}|^{4}{\rm d}s\right]<+\infty$. Therefore we have
@@
+$\quad\quad+\left\{\lambda\,\frac{\kappa}{2}+\sigma^{2}\left(1+\delta\right)\kappa^{2}\right\}\,|\mathfrak{m}_{\alpha}(\rho_{t})|^{2}+\frac{\sigma^{2}}{2}\delta(\delta+1)d.$
@@
(2.5) $\gamma:=\lambda\left(1-\frac{\kappa}{2}\right)-\sigma^{2}\left(1+\delta\right)-b_{2}\left(\lambda\,\frac{\kappa}{2}+\sigma^{2}\left(1+\delta\right)\kappa^{2}\right)$
$\text{and }\quad C_{2}:=\left(\lambda\,\frac{\kappa}{2}+\sigma^{2}\left(1+\delta\right)\kappa^{2}\right)b_{1}+\frac{\sigma^{2}}{2}\delta(\delta+1)d.$
@@
A configuration of the parameters where (2.3) is guaranteed can be obtained as follows. Recalling the notation in (2.5)-(2.6), we can choose $\delta=1$, then we need to satisfy the inequality
@@
[…]
@@
Conclusion: to guarantee $\gamma>0$, it suffices to choose for example
$\delta=1,\quad\kappa<\frac{1}{2}(1+b_{2})^{-1},\quad\text{ and }\quad\lambda>16\,\sigma^{2}.$
The next result captures the asymptotic behavior when $t\to\infty$ in (2.3), and concludes the first step of the desired global convergence of the rescaled CBO to the global minimum.

### v2

@@ 2. Useful estimates
In what follows, $\|\cdot\|$ denotes the Frobenius norm of a matrix and $|\cdot|$ is the standard Euclidean norm in ${\mathbb{R}}^{d}$; $\mathscr{P}({\mathbb{R}}^{d})$ denotes the space of probability measures on ${\mathbb{R}}^{d}$, and $\mathscr{P}_{p}({\mathbb{R}}^{d})$ with $p\geq 1$ contains all $\mu\in\mathscr{P}({\mathbb{R}}^{d})$ such that $\mu(|\cdot|^{p}):=\int_{{\mathbb{R}}^{d}}|x|^{p}\mu({\rm d}x)<\infty$; it is equipped with pp–Wasserstein distance $\mathbb{W}_{p}(\cdot,\cdot)$.
@@
    $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is bounded from below by $\underline{f}:=\min f$, and there exists some constant $L_{f}>0$ such that
@@
[…]
We note that the bound $f(x)-\underline{f}\leq c_{u}(|x|^{2}+1)$ is an immediate consequence of the Lipschitz growth condition in assumption (1). We retain it here to make the dependence on the constant $c_{u}$ explicit in the subsequent results.
Remark 2.2.
The conditions described in Assumption 2.1 are satisfied in practice by a wide range of functions.
  • Local Lipschitz growth: This condition is quite standard and guarantees, in particular, the well-posedness of the dynamics (the associated SDE).
  • Quadratic growth outside a ball: This assumption is satisfied by essentially any objective function whose minimizer (or minimizers) lie within a known bounded region. In practice, optimization procedures often restrict the search domain to a closed ball. One may therefore truncate the objective function on this ball and extend it with a quadratic term outside, e.g. $f(x)\leftarrow f(x)\mathds{1}_{B}(x)+\bigl(|x|^{2}+C\bigr)\mathds{1}_{B^{c}}(x)$, for some appropriate constant $C$. With this modification, all polynomial objective functions, among many others, fall within our framework.
The mean–field term $\mathfrak{m}_{\alpha}(\mu)$ plays a crucial role in the behavior of the dynamics. Its properties are essential for our analysis to be carried out. Thus we shall recall an estimate on $\mathfrak{m}_{\alpha}(\mu)$ that we borrow from [13, Lemma 3.3].
@@
Let $f(\cdot)$ satisfy Assumption 2.1, then it holds that
@@
(2.2) $b_{1}=M^{2}+b_{2}\quad\text{ and }\quad b_{2}=2\,\frac{c_{u}}{c_{\ell}}\left(1+\frac{1}{\alpha\,M^{2}\,c_{\ell}}\right).$
@@
In fact, as it has been already observed in [13, Remark.1], the constant $b_{2}$ can be chosen to be independent of $\alpha$ for $\alpha\geq 1$.
@@
Let $f(\cdot)$ satisfy Assumption 2.1 and $\rho_{0}\in\mathscr{P}_{p}({\mathbb{R}}^{d})$ for any $p\geq 2$, then the dynamics (1.6) has a unique strong solution and satisfies
(2.3) $\sup_{t\in[0,T]}\,{\mathbb{E}}\left[\,|{\overline{X}}_{t}|^{p}\,\right]<\infty\quad\quad\forall\,0<T<\infty.$
@@
The proof of well–posedness for the particle system (1.7) follows the same reasoning as in [13, Theorem 2.1] or [29, Theorem 2.2], relying on the fact that the coefficients in (1.7) are locally Lipschitz. For the mean-field dynamics (1.6), well-posedness can be established using the Leray-Schauder fixed-point theorem, as demonstrated in [13, Theorem 3.1] or [29, Theorem 2.2]. Finally, the moment estimates in (2.3) are derived from [13, Lemma 3.4] or [29, Lemma 3.5] by using the Burkholder–Davis–Gundy inequality. ∎
@@
Let $f(\cdot)$ satisfy assumption 2.1 and $\rho_{0}\in\mathscr{P}_{4}({\mathbb{R}}^{d})$. Consider $({\overline{X}}_{t})_{t\geq 0}$ with $\rho_{0}$-distributed initial data satisfying (1.6). Then for sufficiently small $\kappa$ there exists some constant $C_{2}$ depending only $\lambda,\sigma,d,\delta,\kappa$ and $f$ such that it holds
(2.4) $\sup_{t\geq 0}\;{\mathbb{E}}\left[\,|{\overline{X}}_{t}|^{2}\,\right]\leq\mathbb{E}\left[\,|{\overline{X}}_{0}|^{2}\,\right]+C_{2}\,.$
Before we proceed with the proof, let us recall Itô’s formula of the sake of self-containedness: denote the dynamics of the process $\overline{X}$ by the shorthand notation
$\mbox{\rm d}\overline{X}_{t}=\bm{\mu}_{t}\,\mbox{\rm d}t+\bm{\varsigma}_{t}\,\mbox{\rm d}B_{t}$
Itô’s formula states that for any function $h\in C^{2}(\mathbb{R}^{d})$ one has
(2.5) $\mbox{\rm d}h(\overline{X}_{t})$ $=\left\{\,\nabla h(\overline{X}_{t})\cdot\bm{\mu}_{t}+\frac{1}{2}\text{trace}(\bm{\varsigma}_{t}\cdot D^{2}h(\overline{X}_{t})\,\bm{\varsigma}_{t})\,\right\}\mbox{\rm d}t$
$\qquad+\nabla h(\overline{X}_{t})\cdot\bm{\varsigma}_{t}\,\mbox{\rm d}B_{t}.$
@@
Let $h(x):=\frac{1}{2}|x|^{2}$, for $x\in\mathbb{R}^{d}$, and let $\overline{X}$ be the process defined in (1.6). Then applying Itô’s formula (2.5) leads to
${\rm d}h({\overline{X}}_{t})$ $=\bigg\{-\lambda\nabla h({\overline{X}}_{t})\cdot\big({\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t})\big)$
@@
This is true in particular because of the well-posedness of the process $t\mapsto{\overline{X}}_{t}$ according to Lemma 2.5 with $X_{0}^{i}\in\mathscr{P}_{4}({\mathbb{R}}^{d})$, and its finite moments which guarantee that $\mathbb{E}\left[\int_{0}^{t}|G(s,{\overline{X}}_{s})|^{2}{\rm d}s\right]\leq C\,{\mathbb{E}}\left[\int_{0}^{t}|{\overline{X}}_{s}|^{4}{\rm d}s\right]<+\infty$. Therefore we have
@@
$\quad\quad+\left\{\lambda\,\frac{\kappa}{2}+\sigma^{2}\left(1+\delta\right)\kappa^{2}\right\}\,|\mathfrak{m}_{\alpha}(\rho_{t})|^{2}+\frac{\sigma^{2}}{2}\delta(\delta+1)d.$
@@
(2.7) $\gamma=\lambda\left(1-\frac{\kappa}{2}\right)-\sigma^{2}\left(1+\delta\right)-b_{2}\left(\lambda\,\frac{\kappa}{2}+\sigma^{2}\left(1+\delta\right)\kappa^{2}\right)$
$\text{and }\quad C_{2}=\left(\lambda\,\frac{\kappa}{2}+\sigma^{2}\left(1+\delta\right)\kappa^{2}\right)b_{1}+\frac{\sigma^{2}}{2}\delta(\delta+1)d.$
@@
An example of a simple configuration of the parameters where (2.4) is guaranteed can be obtained as follows. Recalling the notation in (2.2) and then in (2.7), we can choose for example $\delta=1$, then we need to satisfy the inequality
@@
$\Leftrightarrow\quad$ $\lambda\;>\lambda\,\frac{\kappa}{2}(1+b_{2})+2\,\sigma^{2}(1+b_{2}\kappa^{2})$
$\Leftarrow\quad$ $\lambda\;>\lambda\,\kappa\,(1+b_{2})+4\,\sigma^{2}(1+b_{2}\kappa^{2})$
$\Leftrightarrow\quad$ $\lambda\;(1-\,\kappa\,(1+b_{2}))>4\,\sigma^{2}(1+b_{2}\kappa^{2})$
@@
Conclusion: to guarantee $\gamma>0$, an example of sufficient conditions is when
$\delta=1,\quad\kappa<\frac{1}{2(1+b_{2})},\quad\text{ and }\quad\lambda>16\,\sigma^{2}.$
Lastly, we recall a stability estimate on $\mathfrak{m}_{\alpha}(\mu)$ from [29, Corollary 3.3, Proposition A.3]. It will be used in the proof of the subsequent results.